# A LINAC-based, compact x-ray source by inverse Compton scattering

### v1

@@
A LINAC-based, compact x-ray source by inverse Compton scattering
@@
In this paper we present the first measured x-rays produced with a compact, tunable, inverse Compton scattering-based x-ray source. The experimental flux was $1.2\cdot 10^{3}$ photons per shot, which was comparable to simulations. Additionally, the continuous tunability of photon energy and x-ray polarization control possible with the source are shown. The measured x-ray pulse length was in the picosecond range. By increasing experimental parameters, implementing improvements and further conditioning of the accelerator structure, an expected brilliance of $10^{12}$ photons/(s $\times$ mrad2 $\times$ mm2 $\times$ 0.1% BW) is expected, while maximum photon energies can go to 40 keV. Since the complete electron beamline will fit on a single optical table, it is suitable as in-house x-ray source for e.g. research institutes, companies, museums, and hospitals.
@@ 1 Introduction
The most common x-ray source is the x-ray tube, which has already been put into practice for over a century [17]. The working principle of the tube is by targeting an anode metal with accelerated electrons. Over time, developments such as the use of a rotating or liquid metal anode, increased the performance of the device [3, 18]. The brilliance, a figure-of-merit for the output quality, of contemporary tubes is on the order of $10^{9}-10^{10}$ photons/(s $\times$ mrad2 $\times$ mm2 $\times$ 0.1% BW) [19, 20]. Unfortunately, this is not sufficient for all applications. The demand for even higher brilliance, meaning x-ray beams with increased photon flux, a smaller divergence, and a narrower bandwidth therefore gave rise to new types of x-ray sources: synchrotrons and x-ray free electron lasers (XFELs).
@@
By applying bending magnets, wigglers, and undulators, the accelerator-based facilities reach up to a brilliance of 1025 photons/(s $\times$ mrad2 $\times$ mm2 $\times$ 0.1% BW) [21]. Moreover, the continuous tunability of the photon energy and ultrashort x-ray pulses featured in these sources opens up new research possibilities, such as K-edge subtraction imaging and pump-probe experiments [22, 23, 24]. To achieve this, these facilities are however orders of magnitude larger in size than the most advanced x-ray tube, leading to high construction and maintenance costs. As a result, a combined number of less than fifty synchrotrons and approximately ten XFELs are operational worldwide, limiting the availability for a large group of users [25, 26].
@@
To bridge the gap between demand and accessibility of high brilliance beams, recent advancements in accelerator and laser technologies are exploited to develop compact hard x-ray sources. By replacing a standard undulator with a high-power intense laser pulse, inducing the process of inverse Compton scattering (ICS), the size of a synchrotron-like source can be decreased to laboratory scale. Although the large facilities outperform ICS-based sources in terms of x-ray beam quality, the possibility for an in-house source would remove practical burdens for research institutes, companies and hospitals. Furthermore, unlike conventional x-ray tubes, ICS-based sources are also continuously tunable, generate short x-ray pulses, and offer straighforward polarization control [27]. For this reason, multiple ICS-based x-ray sources are either being developed or in commissioning [28, 29, 30, 31, 32].
@@
Among these ICS projects is Smart*Light, the compact source developed at the Eindhoven University of Technology. Smart*Light incorporates a modified version of an X-band accelerator module of the Compact Linear Collider (CLIC) project, allowing the complete electron beamline to fit on a 3 meter sized optical table. Furthermore, the accelerator structure is adapted to accept 100 keV electron bunches. This permits the use of different types of injectors [33, 34, 35]. The modularity enables the user to tailor the output, which can be matched to the needs of the desired application. Smart*Light potentially improves upon x-ray tubes by an order of magnitude in brilliance and by featuring continuous energy tunability from 5 to 40 keV, while restricting the setup size to a single table allows it to still function as an in-house instrument.
@@
In this paper we present the first x-ray measurements of Smart*Light, demonstrating the viability of a compact ICS source. The paper is organized as follows: First we describe the physical principles behind ICS in section 2. In section 3, we address the philosophy behind the Smart*Light project, expressing what makes Smart*Light unique compared to other ICS sources, and why specific choices were made. In section 4, a short overview of the beamline is introduced. A more detailed design description will be published in a separate publication. Finally, section 5 presents the x-ray results, showing the continuous tunability of the photon energies, the characteristic short and narrow band pulses, and the control in polarization.
@@ 2 Principle of ICS
In essence, ICS is a process where photons gain energy by scattering of relativistic electrons. This up-conversion is manifested through a double Doppler-shift. A configuration to induce ICS is schematically represented in figure 1, where an electron with mass $m_{e}$, normalized velocity $\vec{\beta}=\vec{v}/c$ with corresponding Lorentz factor $\gamma$ and momentum $\vec{p}=m_{e}c\gamma\vec{\beta}$ is incident on a high-intensity laser pulse with wavelength $\lambda_{L}$ and angular frequency $\omega_{L}$. $c$ indicates the velocity of light. In the rest frame of the electron, it experiences a laser pulse with an up-shifted frequency. While in the laser pulse, the electron oscillates at this Doppler shifted frequency, upon which it emits dipole radiation. When transformed back to the lab frame, the additional Doppler-shift increases the frequency of the dipole radiation such that a light pulse with wavelength $\lambda_{\mathrm{X}}$ and angular frequency $\omega_{X}=2\pi c/\lambda_{X}$ is observed. Assuming a collision angle $\theta_{L}$ between the laser and a relativistic beam ($\gamma\gg 1$) while neglecting electron recoil, it can be shown that
@@
where $\theta_{X}$ is the angle between x-ray emission and the propagation direction of the electron [36]. Equation (1) only considers linear interactions, i.e. the dimensionless vector potential of the interaction laser $a_{0}=eE_{0}/(2\pi m_{e}c^{2})\ll 1$, which holds for the experimental parameters regarded throughout this work. Assuming head-on scattering ($\theta_{L}=0$) and Gaussian laser and electron beam profiles, a simple analytic expression for the x-ray yield $N_{X}$ can be derived. For the case when the root-mean-square (rms) pulse length of the electron beam $\tau_{e}$ is smaller than its depth of focus, the electron beam equivalent "Rayleigh length", and the rms pulse length of the laser $\tau_{L}$ is significantly shorter than its Rayleigh length $z_{R}$, i.e. $c\tau_{L}\ll z_{R}$, $N_{X}$ is given by:
@@
In Eq. (2), $\sigma_{T}=8\pi r_{0}^{2}/3$ is the Thomson cross section with $r_{0}\approx 2.82$ fm the classical electron radius, $N_{e}$ is the number of electrons in the electron pulse, $N_{L}$ is the number of photons in the laser pulse, $\sigma_{e}$ the rms electron beam size and $\sigma_{L}$ the rms laser spot size where $\sigma_{L}=w_{0}/2$ with $w_{0}=\sqrt{\lambda_{x}z_{R}/\pi}$ the beam waist. As a result of the Doppler shifts, the x-ray emission will be highly directional along the propagation direction of the electrons. It can be shown that for the depicted configuration in figure 1 with $\theta_{L}=0$ and for small angles of $\theta_{X}$, the scattered angular intensity distribution satisfies
@@
with $\nu$ the angle between the polarization and the $x$-axis, and $\phi$ the angle between the $x$-axis and the direction of the considered emission projected on the $x$-$y$ plane [37]. The latter equation shows that in the plane perpendicular to the polarization, the intensity FWHM lies at $\theta_{X}=1/\gamma$, while in the plane parallel to the polarization the intensity goes to zero at $\theta_{X}=1/\gamma$.
To compare the quality of the x-ray source, the brilliance is often used as figure-of-merit, which is defined as:
@@ 3 The philosophy behind Smart*Light
The Smart*Light project has been set up with an explicit goal in mind: the development of a compact, tunable, high-brilliance x-ray source applicable for experiments and diagnostics conducted in companies, museums, hospitals, and research institutes. From this approach, the three key elements of Smart*Light’s philosophy are derived, forming the guideline upon which decisions were made in the design process. Overall, we opted for a linear accelerator (LINAC) based design because of these key elements. The elements can be summarized as follows:
@@
    A table-top setup enables the possibility for an in-house x-ray source, removing schedule and travel barriers. To construct a beamline of this size, it was specifically chosen to use an X-band RF linear accelerator structure over the more conventional S-band and C-band structures. In comparison to the latter two RF-bands, X-band can reach the highest accelerating gradients and higher repetition rates allowing for a more compact design.
@@
    For users it is of importance that the setup operates on demand without the risk that equipment fails or breaks. For this reason, it was chosen to not push the capability of the system to its limits. An example hereof is the generated acceleration gradients in the X-band structure, which could be trained over 100 MV/m. By staying well below this value, dark current is limited and the chance of detrimental vacuum breakdowns is reduced. Moreover, the user can implement a DC photogun as robust and reliable electron source.
@@
Each of the elements aims to improve the applicability for the user, emphasizing that Smart*Light is developed to be an in-house instrument, similar to the electron microscope. Users will benefit from combination of the synchrotron’s tunability and the x-ray tube’s accessibility.
@@ 4 Smart*Light setup
In this section, each segment of the Smart*Light configuration is discussed separately. A schematic overview of the complete system is shown in figure 2. The values mentioned on e.g. bunch charge and laser pulse energy are values used to obtain the results. Parameters of which the system should be capable of are discussed in section 6. A render of the electron beamline is shown in figure 3.
@@ 4.2 Injector segment
The chosen source of electron bunches is a DC photogun, which is used to create 5 pC bunches with an energy of 100 keV by applying field strengths of 10 MV/m to a copper cathode [33, 42]. The electrons are extracted from the cathode through photo-emission with 266 nm UV laser pulses, which are generated via THG (see section 4.1). Upon creation of the electron bunches, space charge forces causes a rapid expansion in size, which is counteracted in the transverse direction by a focusing solenoid placed directly behind the photogun.
In order to accelerate electron bunches in the X-band LINAC, the 100 keV electron bunch length has to be reduced to the ps level. This is accomplished by a TM010 RF compression cavity, acting as a longitudinal lens [43]. The cavity operates in CW mode at a RF frequency of 1.499275 GHz. The RF phase is actively controlled by a feedback system. The RF signal is the eighth subharmonic of the 11.9942 GHz signal supplied by the master oscillator. Producing electric fields up to 0.5 MV/m, the cavity compresses the electron bunches to approximately 2 ps at the injection point of the accelerator.
@@ 4.3 High-power RF segment
Electrons are accelerated by a traveling wave LINAC, operating in the $32\pi/$ 3-mode at a frequency of 11.9942 GHz. The X-band accelerator is designed specifically to allow injection of 100 keV electron bunches. A more detailed design description will be published in a separate paper. RF pulses supplied to the acceleration structure are generated by a Canon E37123 X-band klystron in combination with a Scandinova K200 modulator. The seed RF signal is derived from the master oscillator. The klystron pulses are compressed and amplified before filling the LINAC by a SLED-I pulse compressor [44]. Ultimately, this combination is capable of producing 500 Hz, 200 ns, 24 MW pulses. In the experiments discussed in this paper, 50 Hz, 200 ns, 6-10 MW average power pulses were used. As a result, accelerating gradients up to 42 MV/m were created, increasing the electron energies from 100 keV up to 21 MeV in less than 0.5 m.
4.4 High energy segment
Before interaction with a laser pulse, the relativistic electron bunch is focused transversely by a solenoid to a small interaction spot in order to increase the x-ray flux, see Eq.(2). After a travel distance of approximately 2.2 m, the bunch arrives at the interaction point (IP), where it collides with an intense femtosecond laser pulse to induce ICS. After interaction, electrons are deflected by a permanent dipole magnet into a beam dump. The generated x-ray pulses pass through the gap of the magnet, after which they are detected. For the results, two different detectors have been used. The x-ray intensity distribution has been measured with a MediPix detector [45], while the x-ray energy has been determined using a hyperspectral x-ray camera [46, 47]. Measurements with the latter detector required an appropriate window to get the x-rays out of the vacuum. For this reason, a 100 µm thick diamond window was installed in the exit port of the interaction chamber. A top view of the interaction chamber is presented in figure 4 .
Figure 4: Schematic top view of the interaction chamber, showing the path taken by the interaction laser (red) and electrons (green) inside the chamber.
@@ 5 Results
The intensity pattern of the first x-rays generated by the Smart*Light setup is presented in 5a. For comparison, the expected signal calculated using a Monte Carlo-based algorithm is depicted in 5b [48]. The measurement was done using the 4 quadrant MediPix detector. The separation between the 4 quadrants is clearly visible in Fig. 5a. Although the detector size was too small to completely capture the x-ray beam, the intensity cone can clearly be seen.

### v2

# First x-rays from a compact and tunable LINAC-based Compton scattering source

@@
First x-rays from a compact and tunable LINAC-based Compton scattering source
@@
In this paper, we present the first measurements of x-rays produced with a compact, narrowband, and tunable inverse Compton scattering-based x-ray source, developed at Eindhoven University of Technology. A flux of $1.2\cdot 10^{3}$ photons per shot was measured, in agreement with simulations. Using a high-resolution spectral camera, we show that the photon energy can be tuned continuously from 5.8 keV to 10.7 keV with a bandwidth of 4%. The measured x-ray pulse length was in the picosecond range. Additionally, we show that the source allows full control over the x-ray polarization control. By optimizing experimental parameters, implementing improvements to the setup and further conditioning of the accelerator structure, a brilliance of $10^{12}$ photons/(s $\times$ mrad2 $\times$ mm2 $\times$ 0.1% BW) can be achieved, with photon energies up to 40 keV. Because the complete electron beamline fits on a single optical table, it is suitable as an in-house x-ray source for university laboratories, industrial production lines, museums, and hospitals.
@@ 1 Introduction
The most common x-ray source is the x-ray tube, which has already been put into practice for over a century [17]. The working principle of the tube is by targeting an anode metal with accelerated electrons. Over time, developments such as the use of a rotating or liquid metal anode, increased the performance of the device [3, 18]. The brilliance, a figure of merit for the output quality, of contemporary tubes is on the order of $10^{9}-10^{10}$ photons/(s $\times$ mrad2 $\times$ mm2 $\times$ 0.1% BW) [19, 20]. Unfortunately, this is not sufficient for all applications. The demand for even higher brilliance, meaning x-ray beams with increased photon flux, a smaller divergence, and a narrower bandwidth gave rise to new types of x-ray sources: synchrotrons and x-ray free electron lasers (XFELs).
@@
By applying bending magnets, wigglers, and undulators, the accelerator-based facilities reach up to a brilliance of 1025 photons/(s $\times$ mrad2 $\times$ mm2 $\times$ 0.1% BW) [21]. Moreover, the continuous tunability of the photon energy and ultrashort x-ray pulses featured in these sources opens up new research possibilities, such as K-edge subtraction imaging and pump-probe experiments [22, 23, 24]. However, to achieve this these facilities are required to be orders of magnitude larger in size than the most advanced x-ray tube, leading to high construction and maintenance costs. As a result, a combined number of less than fifty synchrotrons and approximately ten XFELs are operational worldwide, limiting the availability for a large group of users [25, 26].
@@
To bridge the gap between demand and accessibility of high brilliance beams, recent advancements in accelerator and laser technologies are exploited to develop compact hard x-ray sources. By replacing a standard undulator with a high-power intense laser pulse, inducing the process of inverse Compton scattering (ICS), the size of a synchrotron-like source can be decreased to laboratory scale. Although the large facilities outperform ICS-based sources in terms of x-ray beam quality, the possibility for an in-house source would remove practical burdens for research institutes, companies, and hospitals. Furthermore, unlike conventional x-ray tubes, ICS-based sources are also continuously tunable, generate short x-ray pulses, and offer straighforward polarization control [27]. For this reason, multiple ICS-based x-ray sources are either being developed or in commissioning [28, 29, 30, 31, 32].
@@
Among these ICS projects is Smart*Light, the compact source developed at the Eindhoven University of Technology. Smart*Light incorporates a modified version of an X-band accelerating structure of the Compact Linear Collider (CLIC) project, allowing the complete electron beamline to fit on a three-meter-long optical table. Furthermore, the accelerating structure is adapted to capture 100 keV electron bunches. This permits the use of different types of injectors [33, 34, 35]. The modularity enables the user to tailor the output, which can be matched to the needs of the desired application. Smart*Light potentially improves upon x-ray tubes by an order of magnitude in brilliance and by featuring continuous energy tunability from 5 to 40 keV, while restricting the setup size to a single table allows it to still function as an in-house instrument.
@@
In this paper we present the first x-ray measurements of Smart*Light, demonstrating the viability of the compact ICS source. The paper is organized as follows: First, we describe the physical principles behind ICS in Section 2. In Section 3, we address the philosophy behind the Smart*Light project, expressing what makes Smart*Light unique compared to other ICS sources, and why specific choices were made. In Section 4, a brief overview of the beamline is given. A more detailed description of the complete beamline will be published in a separate publication. Finally, Section 5 presents the x-ray results, showing the continuous tunability of the photon energies, the characteristic short and narrow band pulses, and the control in polarization.
@@ 2 Principle of ICS
In essence, ICS is a process where photons gain energy by scattering off relativistic electrons. This up-conversion is manifested through a double Doppler-shift. A configuration to induce ICS is schematically represented in Fig. 1, where an electron with mass $m_{e}$, normalized velocity $\vec{\beta}=\vec{v}/c$ with corresponding Lorentz factor $\gamma$ and momentum $\vec{p}=m_{e}c\gamma\vec{\beta}$ is incident on a high-intensity laser pulse with wavelength $\lambda_{L}$ and angular frequency $\omega_{L}$, with $c$ the velocity of light. In the rest frame of the electron, it experiences a laser pulse with an up-shifted frequency. While in the laser pulse, the electron oscillates at this Doppler shifted frequency, upon which it emits dipole radiation. When transformed back to the lab frame, the additional Doppler-shift increases the frequency of the dipole radiation such that a light pulse with wavelength $\lambda_{\mathrm{X}}$ and angular frequency $\omega_{X}=2\pi c/\lambda_{X}$ is observed. Assuming a collision angle $\theta_{L}$ between the laser and a relativistic beam ($\gamma\gg 1$) while neglecting electron recoil, it can be shown that
@@
where $\theta_{X}$ is the angle between x-ray emission and the propagation direction of the electron [36]. Eq. (1) only considers linear interactions, i.e., the dimensionless vector potential of the interaction laser $a_{0}=eE_{0}/(2\pi m_{e}c^{2})\ll 1$, which holds for the experimental parameters relevant throughout this work. Assuming head-on scattering ($\theta_{L}=0$) and Gaussian laser and electron beam profiles, a simple analytic expression for the x-ray yield $N_{X}$ can be derived. For the case when the root-mean-square (rms) pulse length of the electron beam $\tau_{e}$ is smaller than its depth of focus, the electron beam equivalent "Rayleigh length", and the rms pulse length of the laser $\tau_{L}$ is significantly shorter than its Rayleigh length $z_{R}$, i.e., $c\tau_{L}\ll z_{R}$, $N_{X}$ is given by:
@@
In Eq. (2), $\sigma_{T}=8\pi r_{0}^{2}/3$ is the Thomson cross-section with $r_{0}\approx 2.82$ fm being the classical electron radius, $N_{e}$ is the number of electrons in the electron pulse, $N_{L}$ is the number of photons in the laser pulse, $\sigma_{e}$ the rms-value of the electron beam’s transverse size and $\sigma_{L}$ the rms laser spot size where $\sigma_{L}=w_{0}/2$ with $w_{0}=\sqrt{\lambda_{x}z_{R}/\pi}$ the beam waist. As a result of the Doppler shifts, the x-ray emission will be highly directional along the propagation direction of the electrons. It can be shown that for the depicted configuration in Fig. 1 with $\theta_{L}=0$ and for small angles of $\theta_{X}$, the scattered angular intensity distribution satisfies
@@
with $\nu$ the angle between the polarization and the $x$-axis, and $\phi$ is the angle between the $x$-axis and the direction of the considered emission projected on the $x$-$y$ plane [37]. The right-hand component shows that in the plane perpendicular to the polarization, the intensity FWHM lies at $\theta_{X}=1/\gamma$, while in the plane parallel to the polarization the intensity goes to zero at $\theta_{X}=1/\gamma$.
To compare the quality of the x-ray source, the brilliance is often used as figure of merit, which is defined as:
@@ 3 The philosophy behind Smart*Light
The Smart*Light project has been set up with an explicit goal in mind: the development of a compact, tunable, high-brilliance x-ray source applicable for experiments and diagnostics conducted in companies, museums, hospitals, and research institutes. From this approach, the three key elements of Smart*Light’s philosophy are derived, forming the guideline upon which decisions were made in the design process. Overall, we opted for a linear accelerator (LINAC) based design to accommodate these key elements, which are detailed below:
@@
    A table-top setup enables the possibility for an in-house x-ray source, removing schedule and travel barriers. To construct a beamline of this size, it was specifically chosen to use an X-band (8-12 GHz) RF linear accelerator structure over the more conventional S-band (2-4 GHz) and C-band (4-8 GHz) structures. In comparison to the latter two RF-bands, X-band can reach the highest accelerating gradients and higher repetition rates allowing for a more compact design.
@@
    For users it is of importance that the setup operates on demand without the risk that equipment fails or breaks. For this reason, it was chosen to not push the capability of the system to its limits. An example hereof is the generated acceleration gradients in the X-band structure, which could be trained over 100 MV/m. By staying well below this value, dark current is limited and the chance of detrimental vacuum breakdowns is reduced. Moreover, the user can implement a DC photogun as a robust and reliable electron source.
@@
Each of the elements aims to improve the applicability for the user, emphasizing that Smart*Light is developed to be an in-house instrument, similar to the electron microscope. Users will benefit from Smart*Light’s capabilities which combine a synchrotron’s tunability with a x-ray tube’s accessibility.
@@ 4 Smart*Light setup
In this section, each segment of the Smart*Light configuration is discussed separately. A schematic overview of the complete system is shown in Fig. 2. The values mentioned on e.g. bunch charge and laser pulse energy are values used to obtain the results. Parameters of which the system should be capable of are discussed in Section 6. A render of the electron beamline is shown in Fig. 3.
@@ 4.2 Injector segment
The electron source is a DC photo-emission electron gun (DC photogun), which is used to create 5 pC bunches with an energy of 100 keV by applying field strengths of 10 MV/m to a copper cathode [33, 42]. The electrons are extracted from the cathode through photo-emission with 266 nm UV laser pulses, which are generated via THG (see Section 4.1). Upon creation of the electron bunches, space charge forces causes a rapid expansion in size, which is counteracted in the transverse direction by a solenoid placed directly behind the photogun.
In order to accelerate electron bunches in the X-band LINAC, the 100 keV electron bunch length has to be reduced to the ps level. This is accomplished by a TM010 RF compression cavity, acting as a longitudinal lens [43]. The cavity operates in CW mode at an RF frequency of 1.499275 GHz. The RF phase is actively controlled by a feedback system. The RF signal is the eighth subharmonic of the 11.9942 GHz signal supplied by the master oscillator. Producing electric fields up to 0.5 MV/m, the cavity compresses the electron bunches to approximately 2 ps at the injection point of the accelerator.
@@ 4.3 High-power RF segment
Electrons are accelerated by a traveling wave LINAC, operating in the $32\pi/$ 3-mode at a frequency of 11.9942 GHz. The X-band accelerator is designed specifically to allow injection of 100 keV electron bunches. A more detailed design description of the accelerator can be found in [44]. RF pulses supplied to the acceleration structure are generated by a Canon E37123 X-band klystron in combination with a Scandinova K200 modulator. The seed RF signal is derived from the master oscillator. The klystron pulses are compressed and amplified before filling the LINAC by a SLED-I pulse compressor [45]. Ultimately, this combination is capable of producing 500 Hz, 200 ns, 24 MW pulses. In the experiments discussed in this paper, 50 Hz, 200 ns, 6-10 MW average power pulses were used. As a result, accelerating gradients up to 42 MV/m were created, increasing the electron energies from 100 keV up to 21 MeV in less than 0.5 m.
4.4 High energy segment and x-ray detection
Before interaction with a laser pulse, the relativistic electron bunch is focused transversely by a solenoid to a small interaction spot in order to increase the x-ray flux, see Eq.(2). After a travel distance of approximately 2.2 m, the bunch arrives at the interaction point (IP), where it collides with an intense femtosecond laser pulse to induce ICS. The laser is focused by an off-axis parabolic mirror. After interaction, electrons are deflected by a permanent dipole magnet into a beam dump. The generated x-ray pulses pass through the gap of the magnet, after which they are detected. For the results, two different detectors have been used. The x-ray intensity distribution has been measured with a MediPix detector [46], while the x-ray energy has been determined using a hyperspectral x-ray camera [47, 48]. For the measurements with the latter detector a 100 µm thick diamond window was installed in the exit port of the interaction chamber to decouple the vacuum of the chamber and the detector. A top view of the interaction chamber is presented in Fig. 4.
Figure 4: Schematic top view of the interaction chamber, showing the path taken by the interaction laser (red) and electrons (green) inside the chamber. The x-rays (purple) pass through the bending magnet, after which they are detected.
@@ 5 Results
The intensity pattern of the first x-rays generated by the Smart*Light setup is presented in Fig. 5a. For comparison, the expected signal calculated using a Monte Carlo-based algorithm is depicted in Fig. 5b [49]. The measurement was done using the 4 quadrant MediPix detector. The separation between the 4 quadrants is clearly visible in Fig. 5a. Although the detector size was too small to completely capture the x-ray beam, the intensity cone can clearly be seen.